# Verifying Device Drivers with Pancake

### v1

@@ 1 Introduction
While there have been a number of prior efforts to verify drivers [Alkassar and Hillebrand, 2008; Alkassar, 2009; Duan and Regehr, 2010; Duan, 2013; Chen et al., 2016; Kim et al., 2008; Penninckx et al., 2012; Möre, 2021], to our knowledge none have yet succeeded on realistic, non-trivial devices, nor have they presented any performance analysis of the drivers verified.
@@
Most of these efforts demanded highly manual interactive theorem proving, sometimes requiring drivers to be written and analysed in assembly. On the other hand, attempts to apply more usable methods like model checking [Kim et al., 2008] and automated deductive verification [Penninckx et al., 2012] left significant gaps between the analysed model and the real code.
@@
Moreover, of the above, only Chen et al. [2016] provided an end-to-end verification story that preserved the driver’s verified correctness from a driver-appropriate systems programming language down to the binary level. Likewise, with no formal semantics or verified compiler, recent proposals like that of [Chen et al., 2024] to verify drivers written in Rust [Klabnik and Nichols, 2017] have no plan to close the semantic gap to the binary.
@@
Devices are commodities: new ones are created all the time. This means that to be practicable, their verification must have a high degree of usability and automation. But verification of C code is made needlessly expensive by C’s complicated semantics. In short, the situation demands a performant systems language with support for usable automated verification, and a means of ensuring the semantics of the verified drivers are preserved down to the binary level.
To this end, we present Pancake, an imperative programming language designed to enable verification of low-level systems code, complete with:
@@
  • a verified compiler from Pancake to binary that leverages the final stages of the verified CakeML compilation stack [Tan et al., 2019]. Section 2 discusses the overall structure of Pancake’s semantics and the compiler, and how the verification of its compiler guarantees that the semantics of drivers and other systems code written in Pancake, once verified, will be preserved down to binary level.
@@
  • an automated deductive verification front-end for Pancake that leverages the Viper verification framework [Müller et al., 2016a], a middle-end for various SMT solver-based verification back-ends. This takes the form of (1) an annotation syntax for Pancake and (2) a transpiler from annotated Pancake to the Viper intermediate language (IL), which we explain in Section 3.
@@
Using this support, we produce the first verification of a device driver for a non-trivial device, a driver for the seL4-based LionsOS [Heiser et al., 2025] for a 1 Gb/s Ethernet card that is used in popular Arm SoCs.
@@
Our experience shows that Pancake’s automated deductive verification support is usable for those with a systems development background, allowing the verification of critical guarantees for practical drivers (Section 4). A first-year PhD student, with a background of mostly systems development, but not formal verification, was able to complete the verification of our Pancake-based Ethernet driver in about three person-months (after manually transcribing a C implementation into Pancake)—a process that is likely significantly faster when applied to further drivers. This demonstrates Pancake as a viable alternative to C for systems-level code, but one with the advantage of accessible end-to-end verification support.
@@
We discuss in Section 6 the trusted computing base (TCB) and threats to the validity of Pancake’s automated deductive verification story and end-to-end semantic preservation guarantees, as well as how we plan to address them with future plans for verification.
@@ 2.1 The Rationale
While C is the de-facto standard systems language, C’s semantics has a number of undesirable properties from a verification standpoint: a complicated memory model, underspecified order of evaluation, and the need to prove the absence of undefined behaviour at almost every step. While the seL4 verification demonstrated that these challenges can be overcome, even when verifying machine code without relying on formal properties of a compiler [Sewell et al., 2013], the cost was high: \$350/SLOC just for verifying the C code [Klein et al., 2009], and this cost continues to impact evolution of the kernel. While using a verified compiler [Leroy, 2009] with the verification toolchain VST [Appel, 2011] can help, this has to date not resulted in verified real-world drivers.
Attempts have been made to achieve better systems programming languages by incorporating advanced language features that make certain safety properties hold by construction. For example, Cogent has a linear type discipline that prevents memory leaks [Amani et al., 2016], Rust’s borrow checker enforces ownership and lifetimes [Klabnik and Nichols, 2017], and Cyclone incorporates garbage collection and ML-style polymorphism [Jim et al., 2002]. Such advanced features can eliminate whole classes of bugs, or at least reduce bug density, but at the cost of making the language semantics and implementation more complicated. Furthermore, garbage collection introduces unpredictable delays that are highly undesirable in low-level systems code. Yet these approaches still fall short of ensuring full functional correctness, and it is unclear how helpful such advanced language features are in achieving it.
Functional correctness proofs produce (and consume) significantly stronger properties than type systems typically guarantee. A stronger type system could give more useful information, but unless it is so powerful (and so undecidable) as to be a full-featured proof calculus, the information we need will almost always be stronger than what the type system provides.
The information provided by a type system is only useful if the type system is sound, but most practical languages have unverified type systems, or type systems with known soundness bugs. Type systems can be verified [Naraschewski and Nipkow, 1999], but type soundness proofs tend to be delicate, and have subtle interactions with seemingly minor changes to a language. Maintaining a type soundness proof for a living language can significantly bog down development.
Moreover, the safety guarantees of a language only hold if no backdoors are used. But in low-level systems programming it is often necessary to break out of a type-safe environment. For example, device driver code must adhere to hardware-specified data locations, layouts and access protocols. Hence driver code written in safe languages must use significant amounts of unsafe code, effectively escapes to C [Astrauskas et al., 2020; Evans et al., 2020], which mostly eliminates the benefit of using a safe language.
Instead of adding more safety features to a language, which tends to make the semantics more complicated, we believe a simple formal semantics will help lead to simple proofs. Of course, such a formal semantics must exist in the first place: despite years of research [Jung et al., 2018; Weiss et al., 2019; Kan et al., 2018; Wang et al., 2018], there is still no complete formal specification of Rust.
@@
Pancake is an unmanaged language with no static type system, at a level of abstraction between C and assembly. The data representation and memory model are kept as simple as possible: the only kinds of data are machine words, code pointers, and structs. Programs cannot inspect the stack, which simplifies semantics. All memory is statically allocated; there is no equivalent of malloc and free. There are no concurrency primitives—making drivers single-threaded [Ryzhyk et al., 2009a, 2010] significantly simplifies verification, maps well onto the modular design of microkernel-based OSes, and is routinely used for drivers on seL4 without undue impact on performance [Heiser et al., 2022, 2025].
2 fun handle_irq()
3{ /@ requires valid_device() @/
4 /@ ensures valid_device() @/
5 var 1 EIR = get_device_EIR();
6 !st32 (REG_BASE + EIR_OFFSET), IRQ_MASK;
@@
8 { /@ invariant valid_device() @/
9 var rx_work = EIR & EIR_RXF_MASK;
10 var tx_work = EIR & EIR_TXF_MASK;
11 if (rx_work) {
12 /@ unfold full_heap_access() @/
13 rx_return();
14 rx_provide();
15 /@ fold full_heap_access() @/
@@
17 if (tx_work) {
18 /@ unfold full_heap_access() @/
19 tx_return();
20 tx_provide();
21 /@ fold full_heap_access() @/
22 }
23 if (!rx_work) {
24 if (!tx_work) { break; }
25 }
26 EIR = get_device_EIR();
27 !st32 (REG_BASE + EIR_OFFSET), IRQ_MASK;
28 }
@@ 26 EIR = get_device_EIR();
2.2 The Language and its Semantics
[…]
Since Pancake is targeting verified low-level code, another key design concern is to give programmers direct access to low-level details without the language getting in their way. This is part of the motivation for Pancake’s perhaps most radical design decision: no static type system, and no distinction between different kinds of data. In sloganeering terms, we might say that Pancake is a language where everything is a machine word. For example, there is no distinction between pointers and integers: it’s all words.
We can choose to treat a word as an integer by adding or subtracting with it, and we can choose to treat a word as a pointer by dereferencing it. This means the programmer is free to do arbitrary pointer arithmetic. This is of course unsafe in general, and we do not attempt to make it safe; rather, we give it a simple and well-defined semantics that can support formal verification, without the need for complicated rules about (say) pointer provenance. In this way, the historically minded reader may notice that the language is rather closer to BCPL than C in terms of design philosophy.
With this in mind, the data representation and memory model of Pancake are kept as simple as possible. There are only three kinds of data: machine words, code pointers, and structs (whose fields are machine words, code pointers, or nested structs). Local variables are stack-allocated, and the language does not allow pointers into the stack. Global data may be stored in a statically allocated global memory region: there is no equivalent of malloc and free.
The operational semantics of Pancake is specified in the same style as CakeML’s: functional big-step semantics [Owens et al., 2016], which uses an evaluation function from programs to results. In standard relational big-step semantics, a program is given meaning by a relation between programs and results. The functional style is similar to a language interpreter, but not necessarily executable. All the intermediate languages used in the compiler have this kind of semantics. This style simplifies formal proofs of compiler correctness by making the semantics more amenable to term rewriting. The semantics of a program is defined in terms of how it communicates with the outside world; specifically, as a possibly infinite trace of I/O events, each of which denotes either a returning foreign function call or a shared memory load/store operation. The semantics is parameterised on a function that models the effects of these observable I/O events, i.e., how they change the state of the outside world and what data they pass back to Pancake.
@@ 2.3 Verified Compiler
Figure 2: Overview of CakeML and Pancake compiler stack.
The Pancake compiler is verified, and thus eliminated from the TCB. This sidesteps the need for fragile validation of the compiler output on a program-by-program basis [Sewell et al., 2013].
We re-use the lower parts of the verified CakeML compiler [Kumar et al., 2014; Tan et al., 2019]. Figure 2 shows the relationship, with CakeML compiler on the right side and the Pancake compiler on the left. The CakeML compiler consists of 36 (term rewriting) passes, indicated by arrows in the diagram, and uses 11 intermediate languages (coloured boxes). It compiles a strict functional language (in the style of Standard ML and OCaml) all the way down to concrete machine code for six target architectures.
CakeML itself is a high-level functional programming language, unsuited for low-level systems programming in resource-constrained environments where predictable performance is important. CakeML’s memory management is all handled by the language runtime, and memory allocation may trigger a stop-the-world garbage collector at any time.
Pancake, in contrast, is explicitly designed to be unmanaged and close to hardware, and, importantly, no runtime. Yet by integration into the CakeML ecosystem, it can reuse many of the existing correctness proofs for the CakeML compiler.
The first few phases of our Pancake compiler goes through two intermediate languages (ILs) that are separate from the CakeML compiler’s ILs. The first compiler phase flattens structs and converts the programs to CrepLang, which is a stepping stone into LoopLang. In LoopLang, we compute minimal live sets and divide loops (including their break and continue statements) into tail-recursive functions that better fit the CakeML IL called WordLang. When the program under compilation is translated from LoopLang to WordLang, all loops are replaced with fast tail-calls, as WordLang has no loops. However, the CakeML compiler is set up to compile tail calls into fast simple jumps in the generated machine code.
@@
Our Pancake compiler has a special feature by which it not only produces the output machine code, but can also predict the concrete maximum stack size that the program can use, as long as the user’s input program did not include any recursive calls that are not in tail position. We have proved that the maximum stack size that the Pancake compiler returns is sufficient for running the Pancake program without running out of stack space. This allows Pancake to guarantee the absence of premature termination arising from stack overflows.
The functional correctness proof of the Pancake compiler follows the style of the CakeML compiler proofs. The top-level Pancake compiler correctness theorem states that the source Pancake program and the compiler-generated machine code exhibit exactly the same semantics, in particular, the same observable behaviour (I/O events), if the compiler-generated code is run with at least as much stack space as its maximum stack requires.
The compiler and the associated proofs are free software, and available online. [^1] http://code.cakeml.org for source, or http://cakeml.org for pre-packaged versions. Note that Pancake is fully integrated into the CakeML compiler, rather than a stand-alone release.
@@ 2.4 Challenges with Device Drivers
CakeML assumes exclusive ownership of its statically allocated memory region and that this memory is not observable by the environment. These assumptions are not valid for memory-mapped device registers, which are special memory locations used for interfacing with device hardware, nor does it hold when devices directly write to memory (DMA).
@@
With this limitation, any interaction with the device from CakeML would have to be mediated by FFI calls to C, adding a layer of indirection as part of the TCB. We eliminate this indirection in Pancake by adding native support for interacting with shared memory.
A key challenge here is that the ISA models used by the compiler backend [Tan et al., 2019] are inherently sequential, and $100\mathord{+}$ KSLOC of compiler correctness proofs inextricably rely on this fact. Previous work has integrated driver models for specific devices directly into an ARM ISA model with an interleaving semantics [Alkassar et al., 2007]. This approach would have to abandon much of the existing proof base, and would not provide the modularity we need for targeting multiple ISAs and devices. We therefore parameterise the language semantics on a model for shared memory, which supports proof reuse and provides flexibility for incorporating arbitrary devices.
We also add shared memory semantics to all compiler ILs from Pancake downwards, and verify compilation all the way down to machine code, from Pancake abstract syntax which we extend with load/store operations dedicated for shared memory regions (ShMemLoad and ShMemStore in Figure 1). These operations represent reads and writes that are treated as observable events, just as in the case of any foreign function calls, while being compiled into the load and store instructions of the target machine-language.
@@
Another challenge arising from the CakeML lineage is the issue of function entry points. Implementing the expected driver interface in Pancake requires a mechanism that allows the functions in question to be called from outside the Pancake program. However, the compilation output inherited from CakeML dictates that execution enters through the main function and exits entirely when the main function returns. Additionally, the main function is preceded by lengthy initialisation, and does not support parameters.
@@
Using this execution flow to implement the driver interface demands workarounds, such as branching in the main function based on data indirectly passed through memory, and manually editing the compiler output to return to the caller. Notably, it also results in a substantial performance penalty from repeated re-initialisation. A calling mechanism with this basis is neither scalable nor performant.
@@
Instead, Pancake’s support for multiple entry points provides a calling mechanism to address these issues. Using this feature, a function specification can be tagged with the export keyword, which extends the generated compiler output to expose the function to the calling conventions of the target platform and restore the state initialised by the main function when the exported function is called. This circumvents the re-initialisation, and handles argument passing and returning to the caller without programmer intervention. It enables calling into the Pancake driver as if it was written in C. The compiler correctness proofs currently account only for the main entry point but not for reentry points.

### v2

@@ 1 Introduction
While there have been prior efforts to verify drivers [Alkassar and Hillebrand, 2008; Alkassar, 2009; Duan and Regehr, 2010; Duan, 2013; Chen et al., 2016; Kim et al., 2008; Penninckx et al., 2012; Möre, 2021], to our knowledge none have yet succeeded on realistic, non-trivial devices, nor have they presented any performance analysis of the drivers verified.
@@
Most of these efforts used highly manual interactive theorem proving, and sometimes required drivers to be written and analysed in assembly. On the other hand, attempts to apply more usable methods like model checking [Kim et al., 2008] and automated deductive verification [Penninckx et al., 2012] left significant gaps between the analysed model and the real code.
@@
Of the above, only Chen et al. [2016] provided an end-to-end verification story that preserved the driver’s correctness from a driver-appropriate systems programming language down to the binary. Likewise, with no formal semantics or verified compiler, recent proposals like that of [Chen et al., 2024] to verify drivers written in Rust [Klabnik and Nichols, 2017] have no plan to close the semantic gap to the binary.
@@
Devices are commodities: new ones are created all the time. Therefore, any practicable driver verification must have a high degree of usability and automation. Drivers are typically written in C, but verification of C code is made needlessly expensive by C’s complicated semantics. In short, the situation demands a performant systems language with support for usable automated verification, and a means of ensuring the properties of verified drivers are preserved at the binary level.
In this paper, we present verification of a performant, real-world Ethernet driver that demonstrates a new approach to efficient development of performant and verified device drivers, achieved through the choice of driver design and the use of Pancake, a new imperative programming language designed specifically for verification of low-level systems code. Namely, we verify the Pancake version of an Ethernet driver for LionsOS [Heiser et al., 2025], which uses single-threaded, single-purpose (“modular”) driver designs as advocated by Ryzhyk et al. [2009a, 2010] and outperforms Linux networking with its standard C driver. The Pancake version of the driver we verify also enjoys the excellent performance owing to the simple design, with reasonable overhead coming from porting to Pancake. It will be made available as a choice of network drivers for LionsOS soon.
Our new programming language Pancake has several features that facilitate verification, the most significant being:
@@
  • a verified compiler from Pancake to binary that leverages the final stages of the verified CakeML compiler [Tan et al., 2019]. This guarantees that the semantics of drivers and other systems code written in Pancake, will be preserved at the binary level. Section 2 discusses the overall structure of Pancake’s semantics and the compiler.
@@
  • an automated deductive verification front-end for Pancake that leverages the Viper verification framework [Müller et al., 2016b], a middle-end for various SMT solver-based verification back-ends. This takes the form of (1) an annotation syntax for Pancake, and (2) a transpiler from annotated Pancake to the Viper intermediate language (IL), which we explain in Section 3.
@@
Using this support, we produce the first verification of a device driver for a non-trivial device, a driver for a popular 1 Gb/s Ethernet card for the seL4-based LionsOS [Heiser et al., 2025]. This paper reports on the design of the Pancake language, mainly on its compiler and semantics, and on the verification of the Pancake Ethernet driver using the Viper front-end for Pancake.
@@
Our experience shows that Pancake’s automated deductive verification support is easy to use for those with a systems development background, allowing the verification of critical guarantees for practical drivers (Section 4). The verification of our Ethernet driver took around three person-months. The driver was ported from C to Pancake and then verified by the one member of the team who had no prior verification experience. We expect the process will be even faster once a library of example verified drivers is available. This demonstrates Pancake as a viable alternative to C for systems-level code, but one with the advantage of accessible end-to-end verification support.
@@
In Section 6 we discuss the trusted computing base (TCB) and threats to the validity of Pancake’s automated deductive verification story and end-to-end semantic preservation, and our plan to address them with further verification.
@@ 2.1 The Rationale
[…]
Many attempts to make better systems programming languages incorporate advanced language features to make safety properties hold by construction. For example, Cogent’s linear type discipline prevents memory leaks [Amani et al., 2016], Rust’s borrow checker enforces ownership and lifetimes [Klabnik and Nichols, 2017], and Cyclone incorporates garbage collection and ML-style polymorphism [Jim et al., 2002]. Such features can eliminate whole classes of bugs, or at least reduce bug density, but at the cost of complicating the language semantics and implementation. Garbage collection introduces unpredictable delays that are highly undesirable in low-level systems code. Yet these approaches still fall short of ensuring full functional correctness, and it is unclear how helpful they are in achieving it.
Functional correctness proofs routinely rely on significantly stronger properties than type systems typically guarantee. A stronger type system can give more properties, but unless it is so powerful (and undecidable) as to be a full-featured proof calculus, functional correctness proofs will need more.
The information provided by a type system is only useful in verification if the type system is sound, but most practical languages have unverified type systems, or type systems with known soundness bugs. Type systems can be verified [Naraschewski and Nipkow, 1999], but type soundness proofs are delicate, and have subtle interactions with even minor language changes. Maintaining a type soundness proof for a living language can significantly bog down development.
Moreover, the safety guarantees of a language only hold if no backdoors are used. But low-level systems programs often need to break out of type-safe environments. For example, driver code must adhere to hardware-specified data locations, layouts and access protocols. Hence driver code in safe languages must use significant amounts of unsafe code, effectively escapes to C [Astrauskas et al., 2020; Evans et al., 2020], which mostly eliminates the benefits of safe languages.
Instead of adding more safety features to a language, which tends to complicate the semantics, we believe simple proofs require a simple formal semantics. Of course, the semantics must exist in the first place: despite years of research [Jung et al., 2018; Weiss et al., 2019; Kan et al., 2018; Wang et al., 2018], there is still no complete formal specification of Rust.
@@
Pancake is an unmanaged language with no static type system, at a level of abstraction between C and assembly. The data representation and memory models are kept as simple as possible. The only kinds of data are machine words, code pointers, and structs. Programs cannot inspect the stack, which simplifies semantics. All memory is statically allocated; there is no equivalent of malloc and free.
2.1.1 Non-Goals
Concurrency
User-level drivers running as separate processes do not need to be multi-threaded. Even for Linux in-kernel drivers, Ryzhyk et al. [2009a, 2010] demonstrate that single-threaded drivers are feasible and performant. They furthermore show that 19% of Linux driver bugs are concurrency bugs, which are automatically eliminated by such a design. Single-threaded drivers are much easier to verify, and map well onto the modular design of microkernel-based OSes. This approach is routinely used for drivers on seL4, including in LionsOS, which outperforms Linux on networking using this design [Heiser et al., 2025].
Therefore, Pancake has no concurrency primitives, and our Ethernet driver has no internal concurrency. We will see that this does not meaningfully inhibit performance.
Rust-level verification
Pancake is not intended to replace Rust; rather, it targets a different niche.
Rust’s safety features benefit developers by allowing them to write more trustworthy code in the absence of verification, and Rust-based verification tools like Verus and Prusti have been very successful. However, Rust is a complex language with no authoritative formal semantics, putting any verification on soft foundations. It also requires trusting a large, unverified compiler, which has been found to yield memory-unsafe binaries from typechecked memory-safe Rust code [Speykious, 2024]. Furthermore, “unsafe” code is required for implementing device drivers in Rust, and existing tools have limited support for verification of unsafe Rust.
This level of verification may be sufficient for certain situations, but we aim for stronger guarantees. We take the compiler out of our TCB by fully verifying it, based on Pancake’s precise semantics. A simpler language with formal semantics makes verification easier in both automated and interactive settings. Overall, what we offer is a framework for producing secure low-level components with a very small TCB, through Pancake’s formal semantics and verified compiler.
2 fun example(1 N, 1 shared_addr) {
3 /@ requires mem_access(N) @/
4 /@ ensures mem_access(N) @/
5 /@ ensures N == old(N) @/
6 /@ ensures shared_addr == old(shared_addr) @/
7 var x = 0;
8 var i = 0;
@@ 2.1.1 Non-Goals
10 /@ invariant 0 <= i && i < N && x <= 42 @/
11 /@ invariant mem_access(N) @/
12 var new_x = lds 1 @base + @biw * i;
13 if (i + 1 == N) {break;}
14 if (new_x > 42) {break;}
15 x = new_x;
16 i = i + 1;
@@
18 /@ assert x <= 42 @/
19 !stw shared_addr, x;
@@
2.2 The Language and its Semantics
Pancake looks and feels like a traditional imperative language (see 1, ignore the /@...@/ annotations for now). Figure 1 shows the current abstract syntax of Pancake. The nuts and bolts should be familiar to any programmer. Mutable variables, if, while—nothing fancy. This is a deliberate design decision: we want Pancake to feel simple and familiar to systems programmers. In our experience, this has been borne out in practice: systems programmers familiar with C have found Pancake easy to learn.
@@ 2.2 The Language and its Semantics
Another key design concern is to give programmers direct access to low-level details without the language getting in their way. This is part of the motivation for Pancake’s perhaps most radical design decision: no static type system, and no distinction between different kinds of data. In sloganeering terms, Pancake is a language where everything is a machine word. For example, there is no distinction between pointers and integers: it’s all words.
We can treat a word as an integer by adding to it, or treat it as a pointer by dereferencing it. The programmer can do pointer arithmetic freely, which is of course unsafe in general. We do not attempt to make it safe; rather, we give it a simple and well-defined semantics that can support formal verification, without the need for complicated rules about (say) pointer provenance. In this way, the historically minded reader may find the language design closer to BCPL than C.
With this in mind, the data representation and memory models of Pancake are kept as simple as possible. There are only three kinds of data: machine words, code pointers, and structs (whose fields are machine words, code pointers, or nested structs). Local variables are stack-allocated, and we do not allow pointers into the stack. Global data may be stored in a statically allocated global memory region.
The operational semantics of Pancake is specified using functional big-step semantics [Owens et al., 2016]. In this style, the core of the semantics is a logical function that looks like an interpreter, but is not necessarily executable. This simplifies formal proofs of compiler correctness by making the semantics more amenable to term rewriting. The semantics of a program is defined in terms of how it communicates with the outside world; specifically, as a possibly infinite trace of I/O events, each of which denotes either a foreign function call or a shared memory load/store operation. The semantics is parameterised on a model of the outside world and the kind of I/O events that can occur.
@@ 2.3 Verified Compiler
Figure 2: CakeML and Pancake compiler stack.
The Pancake compiler is verified, and thus not in the TCB. This sidesteps the need for fragile validation of compiler output on a program-by-program basis [Sewell et al., 2013].
We reuse the lower parts of the verified CakeML compiler [Kumar et al., 2014; Tan et al., 2019], which compiles CakeML programs to machine code for six architectures. Figure 2 shows the CakeML compiler on the right and the Pancake compiler on the left. Arrows denote compilation passes, and yellow boxes denote intermediate representations.
CakeML itself is a high-level functional programming language, unsuited for low-level systems programming. For example, CakeML’s memory management is all handled by the language runtime, and memory allocation may trigger a stop-the-world garbage collection at any time.
By contrast, Pancake is explicitly designed to be unmanaged and close to hardware, and to have no runtime. Yet by integration into the CakeML ecosystem, it can reuse many of the existing correctness proofs for the CakeML compiler. The CakeML compiler provides backends for various target architectures [Tan et al., 2019] verified against detailed models of the target ISA semantics.
The first few phases of the Pancake compiler go through two new intermediate languages (ILs). The first compiler phase flattens structs and converts the programs to CrepLang, which is a stepping stone into LoopLang. In LoopLang, we compute minimal live sets and divide loops (including their break and continue statements) into tail-recursive functions that better fit the CakeML IL called WordLang. When the program under compilation is translated from LoopLang to WordLang, all loops are replaced with fast tail-calls, as WordLang has no loops. Later, as in CakeML, these tail calls will be realised as fast simple jumps in the generated machine code.
@@
Our Pancake compiler has a special feature which produces not only the output machine code, but also predicts the concrete maximum stack size the program can use, as long as the input program has recursive calls only in tail position. We have proved that this bound suffices for running the Pancake program without running out of stack space. This allows Pancake to guarantee the absence of premature termination arising from stack overflows.
The functional correctness proof of the Pancake compiler follows the style of the CakeML proofs. The top-level correctness theorem states that the source Pancake program and the compiler-generated machine code exhibit exactly the same observable I/O events, if the compiler-generated code is run with enough stack space.
The compiler and the associated proofs are open source. [^1] http://code.cakeml.org for source, or http://cakeml.org for pre-packaged versions. Note that Pancake is fully integrated into the CakeML compiler, rather than a stand-alone release.
@@ 2.4 Challenges with Device Drivers
Using CakeML as Pancake’s backend compiler presents a number of issues when compiling device driver code. First, CakeML assumes exclusive ownership of its statically allocated memory region, and that this memory is not observable by the environment. These assumptions are not valid for memory-mapped device registers, which are special memory locations used for interfacing with device hardware, nor do they hold when devices directly write to memory (DMA). They also do not fit the LionsOS model, which heavily depends on shared-memory communication between OS components [Heiser et al., 2025].
@@
With this limitation, any interaction with the device from CakeML would have to be mediated by FFI calls to C, adding a layer of indirection to the TCB. We eliminate this indirection in Pancake by adding native support for shared memory.
[…]
We add shared memory semantics to all compiler ILs from Pancake downwards, and verify the full compilation chain. We give load/store operations on shared memory regions a special syntax (ShMemLoad and ShMemStore in Figure 1). In the semantics, these operations are considered observable events. At the end of the compiler pipeline they are converted to ordinary loads and stores.
@@
Another challenge arising from the CakeML lineage is the issue of function entry points. The CakeML compiler expects code to only ever be called via the main function, and to exit entirely when main returns. Moreover, main is preceded by lengthy initialisation, and does not support parameters.
@@
Device drivers in LionsOS must support several functions as entry points, and for these to support re-entry. Thus, using CakeML’s execution flow demands workarounds, such as branching in the main function based on data indirectly passed through memory, and manually editing the compiler output to return to the caller. Even then, a substantial performance penalty arises from repeated re-initialisation. This is clearly neither scalable nor performant.
@@
We address this by implementing support for multiple entry points. A function declaration can be tagged with the export keyword, which extends the generated compiler output to expose the function to the calling conventions of the target platform, and to restore the state initialised by the main function when the exported function is called. This circumvents the re-initialisation, and handles argument passing and returning to the caller without programmer intervention. It enables calling into the Pancake driver as if it were written in C. The top-level compiler correctness proof currently accounts only for the main entry point in that some of the assumptions are too specific to initialisation, which need to be relaxed to account for the re-entry points.